OpenStudy (anonymous):

Identify the center and intercepts of the conic section (2 points): Center: ( , ) x-intercepts: ( , ) y-intercepts: ( , ) ( , ) ( , ) Domain_________________ range: ______________

OpenStudy (anonymous):

this is the test

OpenStudy (anonymous):

from the figure we can see that center: (0,0) x-intercepts: ( -4,0 ) (4 ,0 ) y-intercepts: ( 0, -2)(0 ,2 )

OpenStudy (anonymous):

domain : -4 <= x <=4 range : -2 <= y <= 2

OpenStudy (anonymous):

@Squirrels

OpenStudy (squirrels):

Sorry I dont know, It seems rizwan knows what she is doing

OpenStudy (anonymous):

do you know anyone that knows this stuff im having a lot of trouble!

OpenStudy (squirrels):

One sec

OpenStudy (squirrels):

@BlackLabel

OpenStudy (anonymous):

Rizwan is correct and he/she determined this by looking at the graph. Glad to help ya cheat on your test :D

OpenStudy (squirrels):

:D

OpenStudy (anonymous):

haha could you click on the attachment and look at the rest of the quiz I don't under stand this stuff

OpenStudy (anonymous):

ehhh not in the mood of spitting answers

OpenStudy (anonymous):

pleaseeee! do you know anyone else to help me !!:(

OpenStudy (anonymous):

@cleats is the right address

OpenStudy (anonymous):

yeah hey @cleats wunna help me (:

OpenStudy (anonymous):

Where can I help? rizwan_uet has the right answer

OpenStudy (anonymous):

this is my quiz I need some help(:

OpenStudy (anonymous):

lol post each question separately

OpenStudy (anonymous):

really! thank you !(:

OpenStudy (anonymous):

I meant openstudy questions but whatever. 2) a) circle b) ellipse (because there's a +) c) parabola d) hyperbola 3) the negative makes the parabola flip over the x-axis so it'd be b

OpenStudy (anonymous):

dunno 4 5) recall the general formula:\[\huge (x-h)^2+(y-k)^2=r^2\] center = (h,k) radius = r (2,-3) 3

OpenStudy (anonymous):

CBF to look at any of the others as I don't remember much about ellipses and hyperbolas

OpenStudy (anonymous):

for 5 what would the radius be

OpenStudy (anonymous):

nevermind i just realized it 3 im dumb!

OpenStudy (anonymous):

that, you are

OpenStudy (anonymous):

thanks your to sweet <3

OpenStudy (anonymous):

you're* It's part of a system I have. I act detached and you try to win over my affections. Works every time.

OpenStudy (anonymous):

hahahaha oo geeze!